Skip to end of metadata
Go to start of metadata




  • Status report from La Serena (Stalder). What data were obtained last night?
  • Data transfer and processing (Gruendl). What was transferred and how long did it take. Processing status. 
  • Quality checks on data processed (MacArthur, Gruendl) Let's discuss issues that arose about the CALIB directories.
  • Plan for night 3 (all). 

Discussion items

  • Initial data test took O(2 hours).
  • Second test... Flats should not show condensation

Data Transfer and processing 


  • Eric on transfer :
    • Still no access to postgress logging db
    • files arriving one before creation  (clock issue)
    • LHN people are looking into the network rate
  • Monika
    • same as last night
    • one dark error - 6 ccds instead of 9
      • to be investigated
      • exp ends ti 47 - Brian says there was an issue with writing
      • so used 9 files not 10
  • General calibraiton areas
    • Robert can not see logs
    • using calibraitons from last week (sitting in calib)
    • change in /projects/shared/comcam - with group write - so write there. Any LSST user can write there though and that's not good for ops.
    • Bob asks for general ops (see
  • rerun last nights with "proper" calibrations (the most recent ones).
  • Bob mentions a report on the rehearsal  especially covering the highlights ... 
    • RG suggest DMTN like last time


Flat most interesting - all have compare to previous night (page 9 obvious huge difference with moisture improved). Should be better with correct bias subtraction. 

Brian Stalder mentions illumination does not have corrector optics so doe snot reach edge of field.

DECAM only allow 0.5% across FOV.

Take next sequence with no change.

Action items